# Evaluating Sparse Autoencoders for Monosemantic Representation

### v1

@@ Introduction
As shown in Figure 2, some neurons exhibit considerable overlap in their activation distributions across concepts, while others show separability. This pattern appears in both the base model and SAEs, suggesting that concept erasure techniques should not treat all activation values in one neuron identically. Instead, we propose a more targeted approach that considers where an activation falls within the distribution. Specifically, values in regions uniquely tied to a concept (i.e., those regions of distribution that are clearly separable from other concepts) should be suppressed more strongly, while those shared across concepts should be dampened more conservatively to preserve other concepts. To achieve this, we introduce Attenuation via Posterior Probabilities (APP), which modulates suppression based on distributional separability.
Attenuation via Posterior Probabilities (APP)
Given all neurons $h^{l}_{j}$ of layer $l$, with individual activations $x^{l}_{j}$ for $j=1,\dots,d$, and a target concept $c_{i}\in C$, our goal is to selectively suppress the activation $x^{l}_{j}$ that is attributable to $c_{i}$, while preserving contributions from other concepts.
We begin by computing the posterior probability that a given activation $x^{l}_{j}$ arose from concept $c_{i}$, under the assumption that all concepts are a priori equally likely:
$\pi_{j,i}(x^{l}_{j})$ $=p\bigl{(}c_{i}\mid h^{l}_{j}=x^{l}_{j}\bigr{)}=\frac{p\bigl{(}h^{l}_{j}=x^{l}_{j}\mid c_{i}\bigr{)}\,p(c_{i})}{\displaystyle\sum_{m=1}^{k}p\bigl{(}h^{l}_{j}=x^{l}_{j}\mid c_{m}\bigr{)}\,p(c_{m})},$
$=\frac{p\bigl{(}h^{l}_{j}=x^{l}_{j}\mid c_{i}\bigr{)}}{\displaystyle\sum_{m=1}^{k}p\bigl{(}h^{l}_{j}=x^{l}_{j}\mid c_{m}\bigr{)}}\;\equiv\;\frac{f_{h^{l}_{j}\mid c_{i}}(x^{l}_{j})}{\displaystyle\sum_{m=1}^{k}f_{h^{l}_{j}\mid c_{m}}(x^{l}_{j})}.$
@@
Baseline Methods
We compare our concept erasure approach with the following baselines:
AURA (Suau et al. 2024): Ranks neurons by AUROC, selects those with $\mathrm{AUROC}>0.5$, and dampens each output $z_{m}$ by $\alpha_{m}$ (0 for perfect classifiers, 1 for random), removing concept signals without hyperparameters.
Range Masking (Haider et al. 2025): Identifies concept-relevant neurons and suppresses activations within their typical range ($\mu\pm 2.5\sigma$), erasing concept-specific signals while preserving others.
Adaptive Dampening (Haider et al. 2025): Identifies concept-relevant neurons and dampens activations in proportion to their distance from the concept mean, (strongly near the mean, minimally farther away), enabling precise concept removal.
Full Masking (Dalvi et al. 2019a; Dai et al. 2022): A simple baseline where all neurons associated with the target concept are completely zeroed out to eliminate the concept from the representation.
Table 2: Concept Erasure Results by Type/Method across datasets (Gemma)
Type Method IMDB AG News Emotions DBpedia SST2
@@
Base APP 0.377 0.194 0.569 0.335 0.257 0.378 0.045 0.089 0.154 0.113 0.104 0.076 0.274 0.121 0.307
Aura -0.038 -0.023 2.419 0.070 0.089 1.789 -0.021 0.0002 1.932 0.092 0.064 2.792 -0.120 -0.035 0.553
Range 0.066 0.033 0.511 0.064 0.087 1.057 0.029 0.100 0.797 0.168 0.109 0.700 0.079 0.030 0.402
Adaptive 0.106 0.048 0.287 0.046 0.082 0.561 0.028 0.081 0.436 0.120 0.099 0.387 0.030 0.023 0.225
Full 0.023 0.012 9.692 0.070 0.070 16.574 0.032 0.107 16.913 0.169 0.102 10.922 0.124 0.030 38.469
SAE width: 65k l0: 93 APP 0.596 0.296 0.140 0.603 0.441 0.232 0.238 0.275 0.058 0.399 0.219 0.114 0.642 0.238 0.057
Aura 0.322 0.208 0.266 0.588 0.365 0.575 0.330 0.299 0.256 0.357 0.213 0.490 0.650 0.176 0.164
Range 0.050 0.018 0.276 0.588 0.121 0.554 0.143 0.023 0.387 0.252 0.056 0.446 0.026 0.007 0.192
Adaptive 0.135 0.066 0.169 0.594 0.209 0.366 0.180 0.080 0.203 0.222 0.083 0.279 0.074 0.029 0.109
Full 0.002 0.0001 2.238 0.606 0.006 4.755 0.118 0.001 5.441 0.275 0.018 3.892 0.000 0.000001 4.030
SAE width: 16k l0: 116 APP 0.484 0.242 0.293 0.584 0.369 0.571 0.267 0.269 0.174 0.312 0.152 0.140 0.446 0.235 0.762
Aura 0.548 0.247 0.420 0.598 0.284 1.252 0.325 0.281 0.450 0.257 0.140 0.836 0.083 0.168 1.375
Range 0.099 0.004 0.454 0.357 0.118 0.886 0.096 0.009 0.445 0.208 0.053 0.446 0.274 0.009 0.839
Adaptive 0.124 0.021 0.308 0.360 0.172 0.680 0.123 0.044 0.291 0.177 0.071 0.288 0.272 0.029 0.778
Full 0.002 0.000005 5.271 0.322 0.006 6.474 0.044 0.0004 6.642 0.187 0.0007 5.639 0.000 0.0000001 5.447
SAE width: 65k l0: 197 APP 0.618 0.305 0.317 0.612 0.447 0.411 0.247 0.283 0.307 0.362 0.235 0.083 0.738 0.295 0.095
Aura 0.397 0.197 0.377 0.611 0.391 0.873 0.304 0.317 0.456 0.297 0.235 0.596 0.767 0.233 0.145
Range 0.026 0.012 0.460 0.294 0.046 0.756 0.128 0.021 0.615 0.292 0.106 0.335 0.005 0.0008 0.148
Adaptive 0.114 0.069 0.345 0.295 0.108 0.552 0.169 0.104 0.466 0.257 0.147 0.205 0.019 0.014 0.109
Full 0.008 0.0007 4.179 0.268 0.0005 7.490 0.086 0.002 6.844 0.216 0.019 6.203 0 0.00002 3.599
SAE width: 16k l0: 285 APP 0.314 0.121 0.658 0.628 0.355 0.877 0.210 0.221 0.424 0.274 0.147 0.167 0.043 0.046 0.628
Aura 0.570 0.239 0.780 0.551 0.340 1.628 0.269 0.293 0.649 0.158 0.111 0.975 0.827 0.201 0.648
Range 0.016 0.004 0.978 0.447 0.064 1.886 0.108 0.011 1.350 0.181 0.057 0.958 0.047 0.0007 0.969
Adaptive 0.027 0.021 0.717 0.429 0.142 1.331 0.117 0.067 0.895 0.145 0.083 0.582 0.023 0.004 0.746
Full 0.00008 0.00005 6.646 0.140 0.0003 16.285 0.056 0.001 8.082 0.153 0.023 9.510 0.009 0.000005 7.479
Metrics
@@
Lastly, to capture the overall impact on the model’s generative ability, we measure the drop in perplexity:
$\mathrm{DPPL}=\mathrm{PPL}_{\mathrm{base}}-\mathrm{PPL}_{\mathrm{post}}.$
Implementation Details
To estimate the densities $f_{h^{l}_{j}\mid c_{i}}(x^{l}_{j})$ required by APP, we use kernel density estimation (KDE). Since naive KDE is computationally expensive, we implement a histogram-based approximation that significantly improves efficiency with minimal impact on accuracy. Full implementation details, including runtime complexity, are provided in Appendix B.

### v2

@@ 4 Concept Erasure
As shown in Figure 2, some neurons exhibit considerable overlap in their activation distributions across concepts, while others show separability. This pattern appears in both the base model and SAEs, suggesting that concept erasure techniques should not treat all activation values in one neuron identically. Instead, we propose a more targeted approach that considers where an activation falls within the distribution. Specifically, values in regions uniquely tied to a concept (i.e., those regions of distribution that are clearly separable from other concepts) should be suppressed more strongly, while regions shared across concepts should be dampened more conservatively to preserve other concepts. To achieve this, we introduce Attenuation via Posterior Probabilities (APP), which modulates suppression based on distributional separability.
4.1 Attenuation via Posterior Probabilities (APP)
Given all neurons $h^{l}_{j}$ of layer $l$, with individual activations $x^{l}_{j}$ for $j=1,\dots,d$, and a target concept $c_{i}\in C$, our goal is to selectively suppress the activation $x^{l}_{j}$ that is attributable to $c_{i}$, while preserving contributions from other concepts. We begin by computing the posterior probability that a given activation $x^{l}_{j}$ arose from concept $c_{i}$, under the assumption that all concepts are a priori equally likely:
$\begin{aligned} \pi_{j,i}(x^{l}_{j})&=p\bigl(c_{i}\mid h^{l}_{j}=x^{l}_{j}\bigr)=\frac{p\bigl(h^{l}_{j}=x^{l}_{j}\mid c_{i}\bigr)\,p(c_{i})}{\displaystyle\sum_{m=1}^{k}p\bigl(h^{l}_{j}=x^{l}_{j}\mid c_{m}\bigr)\,p(c_{m})},\\ &=\frac{p\bigl(h^{l}_{j}=x^{l}_{j}\mid c_{i}\bigr)}{\displaystyle\sum_{m=1}^{k}p\bigl(h^{l}_{j}=x^{l}_{j}\mid c_{m}\bigr)}\;\equiv\;\frac{f_{h^{l}_{j}\mid c_{i}}(x^{l}_{j})}{\displaystyle\sum_{m=1}^{k}f_{h^{l}_{j}\mid c_{m}}(x^{l}_{j})}.\end{aligned}$
@@ 4.1 Attenuation via Posterior Probabilities (APP)
4.2 Baseline Methods
To comprehensively evaluate concept erasure effectiveness, we compare APP (which is a partial suppression method) against three other partial methods and one full-masking baseline.
AURA Suau et al. (2024): Ranks neurons by AUROC, selects those with $\mathrm{AUROC}>0.5$, and dampens their output based on $\mathrm{AUROC}$.
Range Masking Haider et al. (2025): The activations of concept-relevant neurons (highly activated) are suppressed when they fall within their typical range ($\mu\pm 2.5\sigma$).
Adaptive Dampening Haider et al. (2025): The activations of concept-relevant neurons (highly activated) are dampened in proportion to their distance from the concept mean.
Full Masking Dalvi et al. (2019a); Dai et al. (2022); Antverg and Belinkov (2022): concept-relevant neurons (highly activated) are fully zeroed out to eliminate the target concept.
Table 2: Concept Erasure Results by Method and Model Type across Datasets (Gemma-2-2B). Bolded values indicate the best performance, and underlined values denote the second-best. Results are grouped by intervention method (e.g., APP, AURA, Adaptive) and model type (Base vs. SAE variants) across five benchmark datasets.
Type Method POS AG News Emotions DBpedia NER
@@ 4.2 Baseline Methods
Base APP 0.224 0.210 0.571 0.347 0.266 0.394 0.051 0.095 0.151 0.113 0.104 0.076 0.079 0.129 0.201
Aura 0.063 0.091 1.195 0.112 0.092 2.130 -0.019 0.004 2.273 0.092 0.064 2.792 0.033 0.044 1.515
Range 0.151 0.1006 2.258 0.081 0.098 1.026 0.036 0.104 0.784 0.168 0.109 0.700 0.135 0.209 1.978
Adaptive 0.157 0.129 1.158 0.050 0.089 0.545 0.032 0.090 0.428 0.120 0.099 0.387 0.095 0.173 1.052
Full 0.115 0.069 22.873 0.096 0.080 16.914 0.038 0.113 17.438 0.169 0.102 10.922 0.134 0.176 23.598
SAE width: 65k l0: 93 APP 0.302 0.270 0.342 0.601 0.442 0.230 0.231 0.279 0.075 0.399 0.219 0.114 0.507 0.520 0.161
Aura 0.406 0.312 0.360 0.577 0.367 0.592 0.336 0.311 0.286 0.357 0.213 0.490 0.519 0.514 0.359
Range 0.052 0.055 0.741 0.587 0.125 0.542 0.146 0.024 0.413 0.252 0.056 0.446 0.497 0.549 0.433
Adaptive 0.095 0.105 0.533 0.592 0.215 0.358 0.182 0.084 0.226 0.222 0.083 0.279 0.494 0.557 0.308
Full 0 0.0003 1.712 0.602 0.006 4.763 0.120 0.002 5.462 0.275 0.018 3.892 0.382 0.391 3.046
SAE width: 16k l0: 116 APP 0.216 0.191 0.468 0.582 0.369 0.569 0.265 0.274 0.166 0.312 0.152 0.140 0.377 0.397 0.531
Aura 0.250 0.192 0.577 0.596 0.286 1.262 0.339 0.280 0.493 0.257 0.140 0.836 0.389 0.415 0.741
Range 0.061 0.046 0.925 0.505 0.103 0.879 0.099 0.009 0.418 0.208 0.053 0.446 0.298 0.305 0.915
Adaptive 0.097 0.085 0.655 0.508 0.163 0.676 0.128 0.047 0.267 0.177 0.071 0.288 0.329 0.347 0.721
Full 0 0.0005 3.738 0.462 0.001 5.882 0.048 0.0005 6.454 0.187 0.0007 5.639 0.262 0.256 3.872
SAE width: 65k l0: 197 APP 0.350 0.322 0.398 0.615 0.458 0.433 0.247 0.285 0.406 0.362 0.235 0.083 0.530 0.654 0.346
Aura 0.384 0.317 0.576 0.611 0.397 0.940 0.311 0.327 0.584 0.297 0.235 0.596 0.567 0.681 0.510
Range 0.046 0.059 0.902 0.346 0.060 0.748 0.135 0.024 0.729 0.292 0.106 0.335 0.503 0.662 0.613
Adaptive 0.110 0.130 0.634 0.345 0.130 0.554 0.170 0.110 0.574 0.257 0.147 0.205 0.503 0.656 0.815
Full 0 0.00001 5.599 0.344 0.0007 7.422 0.110 0.002 7.613 0.216 0.019 6.203 0.292 0.384 3.625
SAE width: 16k l0: 285 APP 0.382 0.306 0.634 0.628 0.356 0.882 0.212 0.228 0.492 0.274 0.147 0.167 0.423 0.482 0.370
Aura 0.433 0.359 0.778 0.549 0.346 1.671 0.286 0.300 0.788 0.158 0.111 0.975 0.428 0.469 0.718
Range 0.051 0.047 1.356 0.456 0.069 1.852 0.111 0.012 1.395 0.181 0.057 0.958 0.338 0.392 0.924
Adaptive 0.179 0.148 0.975 0.441 0.149 1.312 0.121 0.072 0.948 0.145 0.083 0.582 0.376 0.435 0.669
Full 0 0.001 5.880 0.213 0.0004 16.284 0.062 0.001 8.117 0.153 0.023 9.510 0.167 0.191 8.661
4.3 Metrics
@@
Lastly, to capture the overall impact on the model’s generative ability, we measure the increase in perplexity:
$\mathrm{DPPL}=\mathrm{PPL}_{\mathrm{post}}-\mathrm{PPL}_{\mathrm{base}}.$
Comprehensive implementation details are presented in Appendix D, including our histogram-based KDE for concept-conditioned densities.
4.4 Results and Analysis
Comparison of Intervention Effectiveness: SAEs vs. Base Model. As it can be seen in Table 2, across nearly all settings, we find that partial intervention methods (particularly APP and AURA) consistently achieve higher $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ when applied to SAE representations compared to the base model. Specifically, for APP, SAE-based interventions outperformed the base model in 38 out of 40 comparisons. For AURA, SAE-based interventions were more effective in all 40 cases. In contrast, full masking shows less benefit from SAE representations; 19 out of 40 interventions resulted in better outcomes than when applied to the base model. This discrepancy suggests that coarse suppression methods fail to capitalize on the increased concept separability offered by SAEs. These findings reinforce that SAE representations are more disentangled and that fine-grained, distribution-aware methods are better equipped to exploit this structure for effective concept removal. Furthermore, focusing specifically on the APP intervention (which is a partial intervention), we observe that within the SAE family, increasing capacity consistently enhances intervention quality. In particular, $\Delta_{\mathrm{Conf}}$ consistently increases as we scale from 16 k to 65 k latent dimensions, reflecting improved confidence suppression for the target concept. $\Delta_{\mathrm{Acc}}$ also improves across most datasets, with only minor exceptions, further underscoring the role of latent dimensionality in enabling more precise and effective concept removal.
Partial Interventions Vs Full Interventions. Full masking ranks as the worst‐performing method in 36 out of 40 $\Delta$-metrics on the SAEs and it also produces the largest perplexity increase in all DPPL evaluations across SAEs and the base model. This underscores that distribution-aware partial methods (e.g., APP, AURA), which leverage activation distributions, are far more effective for targeted concept removal than the coarse, distribution-agnostic full-masking approach.
APP is least disruptive and highly competitive on concept erasure. APP consistently achieves the smallest $\mathrm{DPPL}$ across all 25 experiments, making it the least disruptive method. Adaptive follows as the second-best approach, attaining the second-smallest $\mathrm{DPPL}$ in 19 out of 25 cases. For concept removal, APP demonstrates strong and consistent performance; it most frequently achieves the best $\Delta_{Conf}$ (14/25 cases), while ranking first or second in 21 out of 25 settings. Regarding $\Delta_{Acc}$, AURA leads with 14 wins, though APP follows closely and ranks in the top two positions for 23 out of 25 settings. Overall, APP effectively removes target concepts while preserving predictive fluency better than all other baselines.
Comparative Analysis of APP and AURA. The superior effectiveness of AURA and APP on both $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ stems from the fact that they explicitly model not only the target‐concept distribution but also the distributions of all auxiliary concepts. By calibrating their interventions to maximize disruption of $c$ while minimizing collateral effects on $c^{\prime}\neq c$, both methods achieve higher $\Delta$ values than approaches that consider only the target distribution. Between these two, APP pulls ahead of AURA in terms of perplexity ($\mathrm{DPPL}$) because it leverages fine‐grained, activation‐specific damping rather than a single, per‐neuron factor. AURA mutes an “expert” neuron uniformly, regardless of whether a particular activation is highly characteristic of the target concept, whereas APP computes $\pi_{j,i}(x)$ for each activation and suppresses only the portions of the distribution uniquely associated with $c_{i}$. This activation-aware attenuation not only removes the targeted concepts effectively but also best preserves the model’s overall fluency, as evidenced by the smallest $\mathrm{DPPL}$.
Cross-Model Validation (DeepSeek). To verify that our observations are not specific to Gemma-2, we replicate experiments on DeepSeek-R1 (Appendix A). The results show the same trends, SAEs enable more selective concept removal than the base model. Moreover, APP achieves the smallest perplexity degradation ($\mathrm{DPPL}$) across all datasets except one, while remaining highly competitive on the concept-removal metrics ($\Delta_{\mathrm{Acc}}$, $\Delta_{\mathrm{Conf}}$).
4.5 Relation Between Erasure Methods and JS Distance
Figure 3: Separability Score vs. Erasure Ability (Partial)
Figure 4: Separability Score vs. Erasure Ability (Full)
Our distribution-aware separability score $S$ (see Subsection 1) serves as a natural predictor for the precision of concept erasure methods. Intuitively, the more separable a neuron’s activation distributions are, the easier it should be to suppress only the target concept while preserving unrelated behavior. We empirically validate this hypothesis in Figure 3, which plots our JS separability score $S$ (x-axis) against the change in accuracy difference $\Delta\text{Acc}$ (y-axis), aggregating data from both the Gemma and DeepSeek experiments (Both SAE and base models). A strong positive Pearson correlation ($r=0.771$, p < 0.001) emerges for the average performance of partial erasure methods (APP, AURA, Range, and Adaptive), confirming that higher separability reliably predicts more selective accuracy drops on the target concept. Detailed per-method correlations are provided in Appendix F. For comparison, we conducted the same analysis for the full masking approach (Figure 4). While a positive correlation also appears ($r=0.657$, p < 0.001), it is notably weaker than that observed for partial masking. This gap reinforces a key insight that full masking cannot exploit the fine-grained separability of activation distributions, whereas distribution-aware partial methods (e.g., APP, AURA) do.
5 Related Works
5.1 Sparse Autoencoders for Feature Discovery
Sparse Autoencoders (SAEs) have emerged as a powerful method for learning interpretable, monosemantic features from neural network activations Huben et al. (2024). Recent advances have focused on improving reconstruction quality and scaling through architectural and training strategy innovations such as JumpReLU activations, BatchTopK sparsity, gated, and end-to-end training frameworks Rajamanoharan et al. (2024a, b); Gao et al. (2025); Bussmann et al. (2024); Braun et al. (2024). Empirical analyses have validated SAEs’ ability to discover meaningful structures across different domains, from vision-language models to algorithmic patterns like temporal difference learning in LLMs Sun et al. (2025b); Pach et al. (2025); Demircan et al. (2025). Moreover, evaluation studies have highlighted both their utility for interpretability tasks and remaining challenges with polysemantic representations Kantamneni et al. (2025); Minegishi et al. (2025); Karvonen et al. (2025). However, none of these empirical analyses evaluated the separability of activation distributions in SAEs as a measure of polysemanticity.
5.2 SAE-Based Model Control
The interpretable features learned by SAEs enable precise control over language model behavior. Several works have demonstrated effective steering by carefully selecting and manipulating SAE features, with approaches ranging from supervised methods for identifying relevant dimensions to frameworks using hypernetworks Arad et al. (2025); He et al. (2025a, b); Bayat et al. (2025); Sun et al. (2025a); Minegishi et al. (2025). However, prior SAE-based control methods did not utilize posterior probabilities, limiting their precision.
5.3 Base Model Control and Causal Analysis
Complementing SAE-based approaches, researchers have developed techniques for direct activation control and causal analysis in the base language models. General frameworks for transporting activations facilitate intervention across model architectures Rodriguez et al. (2025), while causal tracing methods enable precise localization and editing of specific knowledge or biases Vig et al. (2020); Meng et al. (2022, 2023). These approaches offer foundational tools for probing and manipulating model behavior at the activation level. However, again, they do not leverage posterior probabilities for better intervention and have been applied exclusively to base models, not to representations learned by SAEs.
6 Conclusion
This work presents the first quantitative analysis of monosemanticity in SAEs compared to their dense base models through distributional lens. To better characterize monosemanticity, we introduce an activation distribution-aware concept separability score based on the Jensen–Shannon distance, which captures fine-grained distinctions in neuron activations across concepts. We also demonstrate that SAEs support more precise concept-level interventions than base models, particularly when using partial suppression. Building on this, we propose a new method, Attenuation via Posterior Probabilities, which achieves effective concept removal with least possible side effects.
7 limitations
To make density estimation computationally feasible at scale, APP replaces standard kernel density estimation (KDE) with a histogram-based approximation. While this approach substantially improves efficiency, it also introduces certain limitations. As the number of histogram bins increases, the accuracy of the estimated activation distributions improves, but so does the computational cost. Consequently, achieving the best possible performance of APP, in terms of precise density estimation and separability, requires significantly higher computational resources and runtime. Future work could explore alternative KDE methods to better balance accuracy and efficiency.
Appendix A Deepseek Concept Erasure Experiments
On DeepSeek, our method APP is highly competitive on the concept-removal metrics. Across all 20 comparisons of $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ (5 datasets $\times$ 2 metrics $\times$ Base/SAE), APP ranks as the best or second-best method in 13 cases. By contrast, the next strongest method, Range, achieves a top-two ranking in 12 out of 20 comparisons. This demonstrates that APP performs on par with or better than the existing approaches. Importantly, APP is also the least disruptive method, achieving the lowest increase in perplexity (DPPL) in 8 out of 10 DeepSeek experiments (all except POS for Base and POS for SAE). In one of these two exceptions, APP remains highly competitive with a perplexity increase of 0.269 vs. 0.253 for AURA. Overall, APP combines strong concept-removal performance with the smallest degradation in language modeling quality on DeepSeek.
Additionally, as shown in Table 3, applying interventions in the SAE representation yields larger $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ than in the Base representation across methods: for APP, $\Delta_{\mathrm{Conf}}$ increases on all five datasets and $\Delta_{\mathrm{Acc}}$ increases on 4 out of 5 (slightly lower on Emotions), indicating that SAEs promote greater concept separability and thereby enable more effective concept removal.
Appendix B NER Neurons Activation
As shown in Figure 5, the NER dataset exhibits patterns similar to AG News. Certain neurons in both the base model and SAEs (e.g., leftmost plots) show considerable overlap in their activation distributions across the four classes, indicating limited class discrimination. In contrast, neurons in the middle and right columns reveal more separable activation patterns.
Table 5: Concept Erasure Detailed metrics $D_{\mathrm{Acc}}$, $D^{\prime}_{\mathrm{Acc}}$, $D_{\mathrm{Conf}}$, and $D^{\prime}_{\mathrm{Conf}}$ for DeepSeek-R1.
Type Method POS AG News Emotions DBpedia NER
$D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$ $D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$ $D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$ $D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$ $D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$
Base APP 0.198 0.013 0.199 -0.011 0.193 0.009 0.363 -0.021 0.106 0.007 0.162 -0.029 0.160 0.005 0.242 -0.008 0.095 0.007 0.133 -0.010
Aura 0.081 0.032 0.103 0.034 0.029 0.014 0.125 0.054 0.099 0.032 0.179 0.033 0.241 0.030 0.348 0.067 0.051 0.013 0.074 0.003
Range 0.067 0.011 0.288 0.065 0.303 0.130 0.687 0.469 0.148 0.070 0.315 0.164 0.181 0.032 0.347 0.021 0.144 0.054 0.379 0.123
Adaptive 0.060 0.010 0.140 0.001 0.191 0.069 0.519 0.243 0.121 0.050 0.237 0.077 0.136 0.021 0.228 -0.016 0.121 0.031 0.236 0.039
Full 0.062 0.008 0.291 0.109 0.287 0.171 0.691 0.575 0.147 0.078 0.318 0.183 0.183 0.047 0.353 0.064 0.127 0.060 0.371 0.199
SAE APP 0.325 0.018 0.338 0.011 0.507 0.020 0.674 -0.134 0.098 0.006 0.339 -0.134 0.660 0.008 0.573 -0.085 0.581 0.014 0.838 -0.039
Aura 0.424 0.010 0.455 0.043 0.522 0.008 0.688 -0.086 0.098 0.009 0.339 -0.071 0.664 0.012 0.576 -0.198 0.604 0.015 0.869 -0.014
Range 0.420 0.067 0.479 0.057 0.497 0.044 0.644 -0.202 0.098 0.008 0.329 -0.329 0.515 0.017 0.347 -0.299 0.596 0.055 0.855 0.025
Adaptive 0.375 0.036 0.423 -0.008 0.465 0.032 0.607 -0.224 0.098 0.005 0.328 -0.324 0.499 0.011 0.337 -0.294 0.546 0.045 0.793 0.004
Full 0.422 0.177 0.481 0.174 0.501 0.301 0.652 0.174 0.098 0.008 0.329 -0.415 0.520 0.022 0.350 -0.326 0.603 0.303 0.867 0.378
Appendix C Details of Experiments
The detailed metrics ($D_{\mathrm{Acc}}$, $D^{\prime}_{\mathrm{Acc}}$, $D_{\mathrm{Conf}}$, and $D^{\prime}_{\mathrm{Conf}}$) before subtraction for Gemma-2-2b are reported in Table 4, and those for DeepSeek-R1 are presented in Table 5.
Appendix D Implementation Details